OpenStudy (anonymous):

Simplify the expression: 8(3x - 4). A. 24x - 4 B. 24x + 32 C. 24x - 32

OpenStudy (anonymous):

C

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

all u gotta do is substitite

OpenStudy (anonymous):

C

OpenStudy (diyadiya):

8*3x=24x 8*4=32 so 24x-32 =C

OpenStudy (diyadiya):

:D